OpenStudy (anonymous):

simplify:

OpenStudy (anonymous):

\[5\sqrt{3}(2\sqrt{3}+6\sqrt{19}\)]

OpenStudy (anonymous):

The equation tool didn't display this correctly on my computer. Is your question:\[5\sqrt{3}(2\sqrt{3} + 6\sqrt{19})\] ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Ok, that tool is is still not working on my computer, so I might have to help outside of the tool. The expression is of the form a(b + c). You can "combine" the sqrt(3) factors by changing a(b + c) to ab + ac.

OpenStudy (anonymous):

so distributive property?

OpenStudy (anonymous):

Yes. The sqrt(3) will greatly simplify. For the second term, the one that corresponds to "ac", you can multiply the numbers under the sqrt sign.

OpenStudy (anonymous):

i got 10sqrt2+11sqrt57

OpenStudy (anonymous):

[(m)sqrt(n)][(p)sqrt(q)] = (mp)sqrt(nq) That will help for the second term. For the first term, sqrt(a) x sqrt(a) = a

OpenStudy (anonymous):

you jus confused me badly. lol